Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
mantovani, magazzu - 15:12 Saturday 08 April 2023 (59705) Print this report
Longitudinal wp tests and SRCL filter

This morning the ITF could not be locked (the WE arm cavity can could be barely locked but very noise. This was due to a failure, @ 3:58UTC, of the BS suspension, see Figure 1. Valerio has been contacted and it will detail the issue.

Strange unlock at 8:11UTC due to SRCL corr becoming immediatly large, probably due to a reset of SRCL SET FLT (not understood), figure 2.

The OS has been checked and ideed the 0 value of OS signal corresponds to ~ 0Hz of real optical spring.

Then I've checked the effect of the SRCL set on LSC error signal. Lowering the SRCL set we can slightly improve (really a poor improvement) LSC MICH, no effect at low frequency, but the effect on DARM is dramatic, see Figure 3.

The best choice is then to keep the OS @ 0Hz.

we unlocked for DARM failure at OS @ 15Hz.

I tried also to explore the MICH set to lower the coupling of the RIN. see Figures 4 and 5, a more accurate test has been performed after.

The new corrector for SRCL has been tested (purple new , blue old) the gain is in the 30Hz-60Hz bandwidth and improves the accuracy of a factor 2. Not in the automation. Figure 6.

Then I optimized the fmoderr (of -0.1Hz) and I start again the MICH set study.

The minimum of the RIN at high frequency does not correspond to the balancing og  B1p 56MHz sidebands but on the balancing of the 6MHz. Moreover the low frequency coupling minimum does not correspond to the same position of the high frequency one. see Figure 7.

Minimum of PSTAB 1 & 2 @ MICH 0.6 (B1p 6MHz balanced)

Minumum of SSFS line @ MICH 1.2 (B1p 56MHz balanced)

Minumum of PSTAB 0 @ MICH 2

I tried also to repeat the test of yesterday to lower the modulation index of 56MHz but no effect on the coupling was visibile.

At ~14:45 I've removed the 81MHz to check the effect on the sensitivity. No visible effect was observed.

I repeated the scan of MICH, only to check again the wp. no changes (see figure 8)

I leave 0.65 of MICH, next shift can decide how to proceed.

 

 

Images attached to this report
Comments to this report:
flaminio - 8:19 Sunday 09 April 2023 (59714) Print this report

I looked at the data taken yesterday between 9h00 and 9h30 when the SRCL offset was changed.

The SRCL Input decreased from about 24 to about 14 (top right plot), so a little less than a factor of 2.

The change in the SRCL offset determined an increase in the CMRF (third plot from the left on the top).

The change in the SRCL offset also triggered a worstening of the COMMp alignment (two top-left plots) which is known to cause an increase of the CMRF (see https://logbook.virgo-gw.eu/virgo/?r=59508).

The interferometer unlocked at the end of this test, so probably, the working point was not very stable due to the change in the DARM TF that a change in SRCL offset induced.

It would be good to repeat this test while keeping COMMp aligned and making sure that DARM loop is still doing its job properly.

 

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×